OpenStudy (anonymous):

Please help with Trig-- solve for all radian solutions and x if 0

OpenStudy (amistre64):

2sincos = sin(2x) sin(x) + sin(2x) = 0 or sin(2x) = sin(-x) 2x = -x

OpenStudy (amistre64):

or we could look at it ohter ways... 2sin = -sin/cos 2 sin = tan sin = tan/2

OpenStudy (amistre64):

-tan/2 .... gotta keep them signs

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

What happens now?